4)   Aman is thinking of starting up a personal financial service. The basic office supplies will cost $3250 and the variable cost for each consultation is $10.25. He plans to bill each client $50 per consultation.
a)   What is the break-even quantity of consultations for Aman’s personal financial consultation? (3 marks)
b)   If Aman’s goal is $1000, how many customers would he need to reach his goal? (3 marks)
c)   Find the contribution margin and contribution rate. (3 marks)
Answer #1

a)

Fixed Cost=F=$3250

Variable Cost per consultation=V=$10.25

Price=P=$50 per consultation

Break even Quantity=F/(P-V)=3250/(50-10.25)=81.76 or say 82 consultations

b)

Needed Profit=PR=$1000

Required quantity=(F+PR)/(P-V)=(3250+1000)/(50-10.25)=106.92 or say 107 consultations

c)

Contribution Margin per consultation= P-V=50-10.25=$39.75

Contribution Margin Ratio/Rate=Contribution Margin per consultation/Price=39.75/50=79.50%